OpenStudy (anonymous):

What is the equation of a line that passes through (-3,14) and (1,-14)?

OpenStudy (zpupster):

y=mx+b is the equation of a line m=y2-y1/x2-x1 m is the slope (x,y) ordered pair y2=-14 y1=14 x2= 1 x1=-3 m=y2-y1/x2-x1 plug them in gives--watch your signs here m=-14-(14)/1-(-3) m=?

OpenStudy (anonymous):

that is where i am stuck at

OpenStudy (anonymous):

could you just give me the equation?

OpenStudy (zpupster):

give me the slope and i will help with the equation

OpenStudy (anonymous):

I am a little bit lost on how to get the slope. the question only says what i typed for the question.